OpenStudy (anonymous):

An obtuse triangle with area 12 has two sides of lengths 4 and 10. Find the length of the third side. (There are two answers.) (Use law of cosines)

OpenStudy (anonymous):

I got up to the law of cosines

OpenStudy (anonymous):

do you know the law of cosines?

OpenStudy (anonymous):

yeah I did the law of cosines and I ended up getting 1.45=cos4/5

OpenStudy (anonymous):

you know about arccosine?

OpenStudy (anonymous):

that's where you will get two answers

OpenStudy (anonymous):

nope what is that

OpenStudy (anonymous):

Oh yeah I know this is a two answer problem

OpenStudy (anonymous):

arccosine is the function that undoes cosine

OpenStudy (anonymous):

it's probably on your calculator as \[\cos^{-1} \]

OpenStudy (anonymous):

yeah I have that, but do i divide 1.45 by 4/5 to get cos by itself then take the inverse?

OpenStudy (anonymous):

no, cos (4/5) can't be split like that

OpenStudy (anonymous):

so did I do it wrong?

OpenStudy (anonymous):

let's start from the beginning of the problem.

OpenStudy (anonymous):

okay, so i made a triangle, and I found cos by doing 12=(1/2)(10)(4sinC) and that is how I got cos=+-4/5

OpenStudy (anonymous):

Then I did the law of cosines

OpenStudy (anonymous):

ok... that's confusing. law of cosines has nothing to do with the area of the triangle.

OpenStudy (anonymous):

Well the area of the triangle was given, so I'm trying to find the third side of the triangle idk I'm so confused

OpenStudy (anonymous):

Apparently the overall answers are 13.4 and 7.21

OpenStudy (anonymous):

"okay, so i made a triangle, and I found cos by doing 12=(1/2)(10)(4sinC) and that is how I got cos=+-4/5"

OpenStudy (anonymous):

that step would give you sinC, did you use sin^2 + cos^2 =1 to get cos?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how did you get cos from sin?

OpenStudy (anonymous):

i have no idea i looked at someones paper

OpenStudy (anonymous):

don't do that. it makes things harder sometimes.

OpenStudy (anonymous):

okay so how do I get the answer to this problem?

OpenStudy (anonymous):

you can use sin^2 C + cos^2 C =1 to get cos C

OpenStudy (anonymous):

or you could use arcsine to find angle C

OpenStudy (anonymous):

what would sin=

OpenStudy (anonymous):

12=(1/2)(10)(4sinC)

OpenStudy (anonymous):

12=20 sin C sin C= 12/20=6/10=3/5

OpenStudy (anonymous):

yes i have that

OpenStudy (anonymous):

but where do i go from there??